OpenStudy (anonymous):

!!!!!!!!!!!!!!!!!!!!!!!!PLEASE HELP!!!!!!!!!!!!!!!!!!!!!!!!!!!!! MEDALS AWARDED QUESTION BELOW

OpenStudy (anonymous):

\[Prove: \tan ^{2}\theta +\cos ^{2}\theta=1\]

OpenStudy (anonymous):

I know that it does indeed equal 1 but i cannot figure out how to work it out because im really bad at trig

OpenStudy (anonymous):

@phi can you help?

OpenStudy (anonymous):

@skullpatrol do you know?

OpenStudy (skullpatrol):

nope, I'm stuck :(

OpenStudy (skullpatrol):

are you sure you wrote it down correctly?

OpenStudy (anonymous):

i copied and pasted it so im positive

OpenStudy (anonymous):

wait it is wrong hold on

OpenStudy (anonymous):

\[prove:\tan ^{2}\theta \times \cos ^{2}\theta +\cos ^{2\theta}=1\]

OpenStudy (skullpatrol):

:O

OpenStudy (anonymous):

MY BAD UGH LOL SORRY

OpenStudy (anonymous):

its like i didnt copy and paste the whole thing

OpenStudy (skullpatrol):

close this question and start again please

OpenStudy (anonymous):

okay